# L-equivalence and Fourier–Mukai partners of cubic fourfolds

### v1

@@ Abstract.
We study $\mathbb{L}$-equivalence in the Grothendieck ring of varieties and its interaction with categorical invariants of cubic fourfolds. Assuming a Derived Torelli-type criterion for Kuznetsov components and a mild condition on the discriminant of the transcendental lattice, we prove a counting formula for Fourier–Mukai partners of such cubic fourfolds. As an application, we exhibit cubic fourfolds with a fixed algebraic lattice admitting a unique non-trivial Fourier–Mukai partner and which are trivially $\mathbb{L}$-equivalent. Finally, we show that $\mathbb{L}$-equivalence classes of cubic fourfolds are finite.
@@ 1. Introduction
The class $\mathbb{L}\coloneqq[\mathbb{A}^{1}]\in K_{0}(\operatorname{Var}_{k})$ is called the Lefschetz motive. The Lefschetz motive is a zero-divisor in $K_{0}(\operatorname{Var}_{\mathbb{C}})$, as was proved by Borisov [Bor17]. Following this observation, Kuznetsov and Shinder defined $\mathbb{L}$-equivalence:
Definition 1.1 ([KS18]).
@@
$\mathbb{L}$-equivalence seems to be closely related to derived equivalence, especially for projective hyperkähler manifolds. At the time of writing, many examples of $\mathbb{L}$-equivalent varieties have been constructed. In the case of K3 surfaces, see [KS18, HL18, IMOU20, KKM20] and [SZ20], also related with $\mathbb{L}$-equivalent degree five elliptic curves. In higher dimensions, $\mathbb{L}$-equivalent Calabi-Yau varieties are considered in [Bor17, Martin2016, OR2018, IMOU2019, KR2019, Man2019, BCP2020]. Examples of $\mathbb{L}$-equivalence for hyperkähler manifolds appear in [Oka21, BB2025], for Gushel-Mukai surfaces in [PMS2025], for cubic fourfolds in [FL23].
All currently known examples of $\mathbb{L}$-equivalent hyperkähler manifolds are also derived equivalent. In fact, for very general K3 surfaces, $\mathbb{L}$-equivalence implies derived equivalence [Mei24]. On the other hand, there exist examples of derived equivalent hyperkähler manifolds of $\operatorname{K3}^{[n]}$-type that are not $\mathbb{L}$-equivalent, for any $n\geq 2$ [Mei25]. Notably, there are no known examples of K3 surfaces that are derived equivalent but not $\mathbb{L}$-equivalent.
@@
[Mei25] If $X$ and $Y$ are $\mathbb{L}$-equivalent projective hyperkähler manifolds, then they are derived equivalent.
@@
The component $\mathcal{A}_{X}$ is called the Kuznetsov component of $\mathcal{D}^{b}(X)$. Since cubic fourfolds are Fano varieties (in the sense that their anti-canonical bundles are ample), it follows from Bondal and Orlov’s Reconstruction Theorem [BO01] that any Fourier–Mukai equivalence $\mathcal{D}^{b}(X)\simeq\mathcal{D}^{b}(Y)$ between the derived categories of cubic fourfolds induces an isomorphism $X\simeq Y$. However, there are many examples of non-isomorphic cubic fourfolds $X$, $Y$ whose derived categories have equivalent Kuznetsov components $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$. Such cubic fourfolds are called Fourier–Mukai partners.
We therefore propose the following conjecture related to Conjecture 1.2. This conjecture has also recently appeared in [BB2025].
@@
We note that the $\mathbb{L}$-equivalence relation in the context of cubic fourfolds was first studied by Galkin and Shinder [GS14]. In that paper, the so-called rational defect $\mathcal{M}_{X}$ of a cubic fourfold $X$ plays a crucial role. The rational defect is defined to be
@@
If $X$ is rational, the equation $X]=[\mathbb{P}^{4}]+\mathcal{M}_{X}\cdot\mathbb{L}$ holds in the (non-localised) Grothendieck ring $K_{0}(\operatorname{Var}_{\mathbb{C}})$, and $\mathcal{M}_{X}$ is a sum of smooth projective surfaces [GS14, Lemma 2.1]. In fact, in all of the cases where we can compute $\mathcal{M}_{X}$ for rational cubic fourfolds, we have the equation [KS18, IMOU20]
@@
In this paper, we construct a pair of $\mathbb{L}$-equivalent cubic fourfolds $X$, $Y$ whose classes in the Grothendieck ring are of the form (1.1):
Proposition 1.4 (See Proposition 4.5 for the precise statement).
Suppose $X$ is a cubic fourfold with algebraic lattice $A(X)\subset H^{4}(X,\mathbb{Z})$ given by the lattice
$L_{10}\coloneqq\left(\begin{matrix}3&1&1\\ 1&3&0\\ 1&0&10\end{matrix}\right),$
and suppose that $X$ is very general with this property. Then $X$ has a unique non-trivial Fourier–Mukai partner $Y$, and there exists a unique K3 surface $S$ with the property that
@@
The $\mathbb{L}$-equivalent cubic fourfolds in Proposition 1.4 are trivially $\mathbb{L}$-equivalent, in the sense that the motives $[X]$ and $[Y]$ are equal in $K_{0}(\operatorname{Var}_{\mathbb{C}})$. Other examples of trivially $\mathbb{L}$-equivalent cubic fourfolds can be found in [FL23].
Question 1.5.
@@
This assumption is satisfied for a large class of cubic fourfolds (see Proposition 3.3). Inspired by the works [HLOY02, Ma10, BBM25], we prove the following:
@@
Finally, we study the following equivalence relations for cubic fourfolds (we copy the notation from [BFM24]). For two cubic fourfolds $X$, $Y$, we write:
@@ 6. LF:
The fact that $\mathbb{L}$-equivalent cubic fourfolds are Fourier–Mukai partners was recently independently proved in [BB2025] for very general cubic fourfolds in $\mathcal{C}_{d}$ for $d\equiv 0,2\pmod{6}$ not divisible by $9$. They also prove LE $\implies$ LF and the fact that BE does not imply LE.
For the convenience of the reader, let us complete Diagram 1.2 with the implications that are already known. The implication FM $\implies$ DF was proved recently in [KS25], and the fact that BF $\implies$ DF follows from the proof of the $D$-equivalence conjecture for hyperkähler manifolds of $\operatorname{K3}^{[n]}$-type of [MSYZ25]. The question of whether implication FM $\implies$ BE holds is a conjecture by Huybrechts [Huy2023, Conjecture 7.3.21], and the implication DF $\implies$ BE is conjectured in [BFM24, Conjecture 1.2]. For a survey on these equivalence relations, we refer to [BFM24].
@@ Acknowledgements
We thank Franco Giovenzana for pointing out the paper [FL23] to us, and we would also like to thank Evgeny Shinder for explaining how to prove Proposition 5.8. This work is a contribution to Project A22 in SFB 195 No. 286237555 of DFG. The second author is a member of the INdAM group GNSAGA.
@@ 2.1. Lattice theory
Our main reference for lattice theory is [Nik80]. A lattice is a free, finitely generated abelian group $L$ together with a non-degenerate symmetric integral bilinear form
@@
Let $f\colon N\to L$ be an overlattice, and assume $N$ and $L$ are both even. The following results also have analogues for odd lattices [Nik80]. From (2.1), we obtain a natural chain of finite-index injective group homomorphisms:
@@
[Nik80, Proposition 1.4.2] Let $f\colon N\to L$ and $g\colon N\to L^{\prime}$ be two even overlattices of an even lattice $N$.
@@
[Nik80, Proposition 1.5.1] Let $N$ be an even lattice and let $f\colon N\hookrightarrow L$ and $g\colon N\hookrightarrow L^{\prime}$ be primitive embeddings into even lattices $L$, $L^{\prime}$. Write $T\coloneqq N^{\perp}\subset L$ and $T^{\prime}\coloneqq N^{\perp}\subset L^{\prime}$.
@@
The following lemma was noted in [Huy16, §14.0], and was a crucial part of the technical computations in [Mei25].
@@ 2.2. Cubic fourfolds and K3 surfaces
Our main references for K3 surfaces and cubic fourfolds are [Huy16, Huy2023]. Let $X$ be a cubic fourfold. We consider the cohomology group $H^{4}(X,\mathbb{Z})$. The intersection product defines a lattice structure on $H^{4}(X,\mathbb{Z})$, which abstractly makes it isometric to the lattice
@@
[Voisin1986] Let $X$ and $Y$ be two cubic fourfolds. Then $X$ and $Y$ are isomorphic if and only if there exists a Hodge isometry $H^{4}(X,\mathbb{Z})_{\text{prim}}\simeq H^{4}(Y,\mathbb{Z})_{\text{prim}}$.
@@
A fundamental tool for studying cubic fourfolds is the period map, which associates a cubic fourfold $X$ in its moduli space to its primitive cohomology $H^{4}_{\operatorname{prim}}(X,\mathbb{Z})$, viewed as a point in the corresponding period domain, see [Huy2023] for more details.
@@
[Looijenga2009, Laza2010] The image of the period map for cubic fourfolds is the complement of the divisors $\mathcal{C}_{2}\cup\mathcal{C}_{6}$.
It follows from the surjectivity of the period map that $A(X)_{\mathrm{prim}}$ contains no vectors of square $2$, nor any vectors of square $6$ and divisibility $3$. For this reason, we define:
@@
The derived category of a cubic fourfold $X$ admits a semiorthogonal decomposition [Kuz2010]:
@@
The component $\mathcal{A}_{X}$ is a K3 category called the Kuznetsov component. The topological Grothendieck group $K_{\text{top}}(\mathcal{A}_{X})$ has a natural Hodge structure of K3-type of weight $2$ as defined in [AT2014]. The resulting Hodge lattice is usually denoted $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$. As a lattice, $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ is isometric to the even, unimodular lattice of signature $(20,4)$:
@@
[Muk87, Orl03] Let $S$ and $S^{\prime}$ be K3 surfaces. The following are equivalent:
@@
Returning to cubic fourfolds, it is known for several families of special cubic fourfolds that the Kuznetsov component is explicitly realised as the derived category of a (possibly twisted) K3 surface: for cubic fourfolds containing a plane this goes back to [Kuz2010] and has been extended to the non-generic nodal case in [Mos2018]. Recently this has been studied for cubic fourfolds containing multiple planes in [Hart2025]. However, it is currently an open conjecture that Theorem 2.12 has an analogue for the Kuznetsov components of cubic fourfolds. That is, for two cubic fourfolds $X,Y$, it is currently not known whether there is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ if there is a Hodge isometry $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})\simeq\tilde{H}(\mathcal{A}_{Y},\mathbb{Z})$. There is the following positive result, however:
@@
[HS05, AT2014] Let $X$ and $Y$ be cubic fourfolds. Consider the following three statements:
@@ 2.3. Hodge realisation
We briefly recall the constructions of [Efi18, Mei24, Mei25]. We denote by $\operatorname{HS}_{\mathbb{Z}}$ the category of pure, graded, polarisable, integral Hodge structures. The Grothendieck group of Hodge structures is defined to be the group
@@
[Mei24, Mei25] There is a group homomorphism
@@
Addington showed in [Add16, Corollary 8] that there is a natural embedding $A_{2}=\langle\lambda_{1},\lambda_{2}\rangle\subset\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ such that
@@
where the (1)(-1) indicates that we rescaled the bilinear forms by 1-1. Since the divisibility of $\lambda_{1}$ in $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ is $1$, it follows that the coarseness of $F(X)$ is $\operatorname{crs}(F(X))=1$ (see [Mei25, Definition 2.4] for the definition of coarseness). Therefore, by [Mei25, Theorem 3.5], we find
@@
Suppose that there is a K3 surface $S$ with a Hodge isometry $T(X)(-1)\simeq T(S)$, so that $F(X)$ is a moduli space of stable sheaves on $S$ by [Bec22, Corollary 9.6]. Then Lemma 2.19 is equivalent to saying that $F(X)$ is a fine moduli space of sheaves on $S$ by [MM24, Theorem 3.9].
The following result appears as Lemma 3.2 in [BB2025].
@@
Let $X$ and $Y$ be cubic fourfolds (resp. K3 surfaces). Suppose $\operatorname{rk}(T(X))\neq 4$, and that $\operatorname{End}_{\text{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, if $[T(X)]=[T(Y)]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z}})$, there is a Hodge isometry $T(X)\simeq T(Y)$.
@@
If $X$ and $Y$ are K3 surfaces, the statement is follows from [Mei24, Proposition 3.11] combined with [Efi18, Lemma 3.7].
If $X$ and $Y$ are cubic fourfolds, we have $[T(X)(1)]=[T(Y)(1)]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},2})$. Since $\operatorname{End}_{\text{Hodge}}(T(X))\simeq\mathbb{Z}$, this implies by [Efi18, Lemma 3.7] that there exists an isomorphism of Hodge structures (not necessarily a Hodge isometry) $T(X)(1)\simeq T(Y)(1)$. Moreover, since $X$ and $Y$ are $\mathbb{L}$-equivalent, we have $\operatorname{disc}(T(X))=\operatorname{disc}(T(Y))$ by Lemma 2.17 and Lemma 2.18. It now follows from [Mei24, Corollary 3.10] that we have a Hodge isometry $T(X)(1)\simeq T(Y)(1)$, which clearly induces a Hodge isometry $T(X)\simeq T(Y)$ via the Tate twist. ∎

### v2

@@ Abstract.
We study $\mathbb{L}$-equivalence in the Grothendieck ring of varieties and its interaction with categorical invariants of cubic fourfolds. Assuming a Derived Torelli-type criterion for Kuznetsov components and a mild condition on the discriminant of the transcendental lattice, we prove a counting formula for Fourier–Mukai partners of such cubic fourfolds. As an application, we exhibit cubic fourfolds with a fixed algebraic lattice admitting a unique non-trivial Fourier–Mukai partner, which is trivially $\mathbb{L}$-equivalent to the original. Finally, we show that $\mathbb{L}$-equivalence classes of cubic fourfolds are finite.
@@ 1. Introduction
The class $\mathbb{L}\coloneqq[\mathbb{A}^{1}]\in K_{0}(\operatorname{Var}_{k})$ is called the Lefschetz motive. The Lefschetz motive is a zero-divisor in $K_{0}(\operatorname{Var}_{\mathbb{C}})$, as was proved by Borisov [6]. Following this observation, Kuznetsov and Shinder defined $\mathbb{L}$-equivalence:
Definition 1.1 ([30]).
@@
$\mathbb{L}$-equivalence seems to be closely related to derived equivalence, especially for projective hyperkähler manifolds. At the time of writing, many examples of $\mathbb{L}$-equivalent varieties have been constructed. In the case of K3 surfaces, see [30, 18, 25, 27] and [52], also related with $\mathbb{L}$-equivalent degree five elliptic curves. In higher dimensions, $\mathbb{L}$-equivalent Calabi-Yau varieties are considered in [6, 36, 49, 26, 28, 35, 7]. Examples of $\mathbb{L}$-equivalence for hyperkähler manifolds appear in [46, 4], for Gushel-Mukai surfaces in [51], for cubic fourfolds in [13].
All currently known examples of $\mathbb{L}$-equivalent hyperkähler manifolds are also derived equivalent. In fact, for very general K3 surfaces, $\mathbb{L}$-equivalence implies derived equivalence [39]. On the other hand, there exist examples of derived equivalent hyperkähler manifolds of $\operatorname{K3}^{[n]}$-type that are not $\mathbb{L}$-equivalent, for any $n\geq 2$ [40]. Notably, there are no known examples of K3 surfaces that are derived equivalent but not $\mathbb{L}$-equivalent.
@@
[40] If $X$ and $Y$ are $\mathbb{L}$-equivalent projective hyperkähler manifolds, then they are derived equivalent.
@@
The component $\mathcal{A}_{X}$ is called the Kuznetsov component of $\mathcal{D}^{b}(X)$. Since cubic fourfolds are Fano varieties (in the sense that their anti-canonical bundles are ample), it follows from Bondal and Orlov’s Reconstruction Theorem [5] that any Fourier–Mukai equivalence $\mathcal{D}^{b}(X)\simeq\mathcal{D}^{b}(Y)$ between the derived categories of cubic fourfolds induces an isomorphism $X\simeq Y$. However, there are many examples of non-isomorphic cubic fourfolds $X$, $Y$ whose derived categories have equivalent Kuznetsov components $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$. Such cubic fourfolds are called Fourier–Mukai partners.
We therefore propose the following conjecture related to Conjecture 1.2. This conjecture has also recently appeared in [4].
@@
Remark 1.4.
Conjecture 1.3 should also follow from Kontsevich’s expectation that semiorthogonal decompositions are canonical (up to mutation) and functorial with respect to geometric operations. See, for further references [12, 15].
We note that the $\mathbb{L}$-equivalence relation in the context of cubic fourfolds was first studied by Galkin and Shinder [14]. In that paper, the so-called rational defect $\mathcal{M}_{X}$ of a cubic fourfold $X$ plays a crucial role. The rational defect is defined to be
@@
If $X$ is rational, the equation $X]=[\mathbb{P}^{4}]+\mathcal{M}_{X}\cdot\mathbb{L}$ holds in the (non-localised) Grothendieck ring $K_{0}(\operatorname{Var}_{\mathbb{C}})$, and $\mathcal{M}_{X}$ is a sum of smooth projective surfaces [14, Lemma 2.1]. In fact, in all of the cases where we can compute $\mathcal{M}_{X}$ for rational cubic fourfolds, we have the equation [30, 25]
@@
In this paper, we construct pairs of $\mathbb{L}$-equivalent cubic fourfolds $X$, $Y$ whose classes in the Grothendieck ring are of the form (1.1):
Proposition 1.5 (See Proposition 4.5 and Proposition 4.7).
Suppose $X$ is a cubic fourfold with one of the following two properties:
  1. A)
    The primitive part $A(X)_{\mathrm{prim}}$ of the algebraic lattice is isometric to
    $L_{2,9}\coloneqq\left(\begin{matrix}14&2\\ 2&18\end{matrix}\right).$
  2. B)
    The algebraic lattice $A(X)\subset H^{4}(X,\mathbb{Z})$ is isometric to the lattice
    $L_{10}\coloneqq\left(\begin{matrix}3&1&1\\ 1&3&0\\ 1&0&10\end{matrix}\right).$
Moreover, suppose that $X$ is very general with one of these properties. Then $X$ has a non-trivial Fourier–Mukai partner $Y$ for which there exists a unique K3 surface $S$ with the property that
@@
Finally, if $X$ satisfies Property B, then $Y$ is the unique non-trivial Fourier–Mukai partner of $X$.
The $\mathbb{L}$-equivalent cubic fourfolds in Proposition 1.5 are trivially $\mathbb{L}$-equivalent, in the sense that the motives $[X]$ and $[Y]$ are equal in $K_{0}(\operatorname{Var}_{\mathbb{C}})$. Other examples of trivially $\mathbb{L}$-equivalent cubic fourfolds can be found in [13].
Question 1.6.
@@
This assumption is satisfied for a large class of cubic fourfolds (see Proposition 3.3). Inspired by the works [19, 34, 9], we prove the following:
@@
We also prove a partial counting formula for Fourier–Mukai partners of a cubic fourfold $X$ of which the primitive algebraic lattice $N\coloneqq A(X)_{\mathrm{prim}}$ has discriminant indivisible by $3$. For such a cubic fourfold, we prove in Proposition 3.12 that the number of Fourier–Mukai partners $Y$ which satisfy $A(Y)_{\mathrm{prim}}\simeq N^{\prime}\in\mathcal{H}(N)$ is given by
$|O(N^{\prime})\backslash O(A_{N})/O_{\mathrm{Hodge}}(T(X))|.$
Note that $X$ may have Fourier–Mukai partners $Y$ for which the discriminant of $A(Y)_{\mathrm{prim}}$ is a multiple of $3$. In this case we have $A(Y)_{\mathrm{prim}}\notin\mathcal{H}(N)$. We give examples of this phenomenon in Remark 4.8.
Finally, we study the following equivalence relations for cubic fourfolds (we copy the notation from [8]). For two cubic fourfolds $X$, $Y$, we write:
@@ 6. LF:
The fact that $\mathbb{L}$-equivalent cubic fourfolds are Fourier–Mukai partners was recently independently proved in [4] for very general cubic fourfolds in $\mathcal{C}_{d}$ for $d\equiv 0,2\pmod{6}$ not divisible by $9$. We note that such cubic fourfolds satisfy Assumption 1.7. They also prove LE $\implies$ LF and the fact that BE does not imply LE.
For the convenience of the reader, let us complete Diagram 1.2 with the implications that are already known. The implication FM $\implies$ DF was proved recently in [29], and the fact that BF $\implies$ DF follows from the proof of the $D$-equivalence conjecture for hyperkähler manifolds of $\operatorname{K3}^{[n]}$-type of [37]. Finally, under Assumption 1.7, we also have DF $\implies$ FM by [3, Corollary 9.3]. The question of whether implication FM $\implies$ BE holds is a conjecture by Huybrechts [24, Conjecture 7.3.21], and the implication BF $\implies$ BE is conjectured in [8, Conjecture 1.2]. For a survey on these equivalence relations, we refer to [8].
@@ Acknowledgements
We thank Franco Giovenzana for pointing out the paper [13] to us, and we would also like to thank Evgeny Shinder for telling us about Proposition 5.8. Finally, we thank Corey Brooke, Sarah Frei, and Lisa Marquand for useful comments on an earlier version of this paper. This work is a contribution to Project A22 in SFB 195 No. 286237555 of DFG. The second author is a member of the INdAM group GNSAGA.
@@ 2.1. Lattice theory
Our main reference for lattice theory is [44]. A lattice is a free, finitely generated abelian group $L$ together with a non-degenerate symmetric integral bilinear form
@@
Let $f\colon N\to L$ be an overlattice, and assume $N$ and $L$ are both even. The following results also have analogues for odd lattices [44]. From (2.1), we obtain a natural chain of finite-index injective group homomorphisms:
@@
[44, Proposition 1.4.2] Let $f\colon N\to L$ and $g\colon N\to L^{\prime}$ be two even overlattices of an even lattice $N$.
@@
[44, Proposition 1.5.1] Let $N$ be an even lattice and let $f\colon N\hookrightarrow L$ and $g\colon N\hookrightarrow L^{\prime}$ be primitive embeddings into even lattices $L$, $L^{\prime}$. Write $T\coloneqq N^{\perp}\subset L$ and $T^{\prime}\coloneqq N^{\perp}\subset L^{\prime}$.
@@
The following lemma was noted in [21, §14.0], and was a crucial part of the technical computations in [40].
@@ 2.2. Cubic fourfolds and K3 surfaces
Our main references for K3 surfaces and cubic fourfolds are [21, 24]. Let $X$ be a cubic fourfold. We consider the cohomology group $H^{4}(X,\mathbb{Z})$. The intersection product defines a lattice structure on $H^{4}(X,\mathbb{Z})$, which abstractly makes it isometric to the lattice
@@
[53] Let $X$ and $Y$ be two cubic fourfolds. Then $X$ and $Y$ are isomorphic if and only if there exists a Hodge isometry $H^{4}(X,\mathbb{Z})_{\text{prim}}\simeq H^{4}(Y,\mathbb{Z})_{\text{prim}}$.
@@
A fundamental tool for studying cubic fourfolds is the period map, which associates a cubic fourfold $X$ in its moduli space to its primitive cohomology $H^{4}_{\operatorname{prim}}(X,\mathbb{Z})$, viewed as a point in the corresponding period domain, see [24] for more details.
@@
[33, 32] The image of the period map for cubic fourfolds is the complement of the divisors $\mathcal{C}_{2}\cup\mathcal{C}_{6}$.
It follows from the surjectivity of the period map that $A(X)_{\mathrm{prim}}$ contains no vectors of square $2$, nor any vectors of square $6$ and divisibility $3$ in $H^{4}(X,\mathbb{Z})_{\mathrm{prim}}$. For this reason, we define:
@@
The derived category of a cubic fourfold $X$ admits a semiorthogonal decomposition [31]:
@@
The component $\mathcal{A}_{X}$ is a K3 category called the Kuznetsov component. The topological Grothendieck group $K_{\text{top}}(\mathcal{A}_{X})$ has a natural Hodge structure of K3-type of weight $2$ as defined in [2]. The resulting Hodge lattice is usually denoted $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$. As a lattice, $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ is isometric to the even, unimodular lattice of signature $(20,4)$:
@@
[42, 47] Let $S$ and $S^{\prime}$ be K3 surfaces. The following are equivalent:
@@
Returning to cubic fourfolds, it is known for several families of special cubic fourfolds that the Kuznetsov component is explicitly realised as the derived category of a (possibly twisted) K3 surface: for cubic fourfolds containing a plane this goes back to [31] and has been extended to the non-generic nodal case in [41]. Recently this has been studied for cubic fourfolds containing multiple planes in [17]. However, it is currently an open conjecture that Theorem 2.12 has an analogue for the Kuznetsov components of cubic fourfolds. That is, for two cubic fourfolds $X,Y$, it is currently not known whether there is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ if there is a Hodge isometry $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})\simeq\tilde{H}(\mathcal{A}_{Y},\mathbb{Z})$. There is the following positive result, however:
@@
[20, 2] Let $X$ and $Y$ be cubic fourfolds. Consider the following three statements:
@@ 2.3. Hodge realisation
We briefly recall the constructions of [11, 39, 40]. We denote by $\operatorname{HS}_{\mathbb{Z}}$ the category of pure, graded, polarisable, integral Hodge structures. The Grothendieck group of Hodge structures is defined to be the group
@@
[39, 40] There is a group homomorphism
@@
Addington showed in [1, Corollary 8] that there is a natural embedding $A_{2}=\langle\lambda_{1},\lambda_{2}\rangle\subset\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ such that
@@
where the (1)(-1) indicates that we rescaled the bilinear forms by 1-1. Since the divisibility of $\lambda_{1}$ in $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ is $1$, it follows that the coarseness of $F(X)$ is $\operatorname{crs}(F(X))=1$ (see [40, Definition 2.4] for the definition of coarseness). Therefore, by [40, Theorem 3.5], we find
@@
Suppose that there is a K3 surface $S$ with a Hodge isometry $T(X)(-1)\simeq T(S)$, so that $F(X)$ is a moduli space of stable sheaves on $S$ by [3, Corollary 9.6]. Then Lemma 2.19 is equivalent to saying that $F(X)$ is a fine moduli space of sheaves on $S$ by [38, Theorem 3.9].
The following result appears as Lemma 3.2 in [4].
@@
Let $X$ and $Y$ be cubic fourfolds (resp. K3 surfaces). Suppose $\operatorname{rk}(T(X))\neq 4$, and that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, if $[T(X)]=[T(Y)]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z}})$, there is a Hodge isometry $T(X)\simeq T(Y)$.
@@
If $X$ and $Y$ are K3 surfaces, the statement is follows from [39, Proposition 3.11] combined with [11, Lemma 3.7].
If $X$ and $Y$ are cubic fourfolds, we have $[T(X)(1)]=[T(Y)(1)]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},2})$. Since $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$, this implies by [11, Lemma 3.7] that there exists an isomorphism of Hodge structures (not necessarily a Hodge isometry) $T(X)(1)\simeq T(Y)(1)$. Moreover, since $X$ and $Y$ are $\mathbb{L}$-equivalent, we have $\operatorname{disc}(T(X))=\operatorname{disc}(T(Y))$ by Lemma 2.17 and Lemma 2.18. It now follows from [39, Corollary 3.10] that we have a Hodge isometry $T(X)(1)\simeq T(Y)(1)$, which clearly induces a Hodge isometry $T(X)\simeq T(Y)$ via the Tate twist. ∎